# Inflationary Gravitational Wave Spectral Shapes as test for Low-Scale Leptogenesis

### v1

@@ Abstract
We study thermal and non-thermal resonant leptogenesis in a general setting where a heavy scalar $\phi$ decays to right-handed neutrinos (RHNs) whose further out-of-equilibrium decay generates the required lepton asymmetry. Domination of the energy budget of the Universe by the $\phi$ or the RHNs alters the evolution history of the primordial gravitational waves (PGW), of inflationary origin, which re-enter the horizon after inflation, modifying the spectral shape. The decays of $\phi$ and RHNs release entropy into the early Universe while nearly degenerate RHNs facilitate low and intermediate scale leptogenesis. We show that depending on the coupling $y_{R}$ of $\phi$ to radiation species, RHNs can achieve thermal abundance before decaying, which gives rise to thermal leptogenesis. A characteristic damping of the GW spectrum resulting in two knee-like features or one knee-like feature would provide evidence for low-scale thermal and non-thermal leptogenesis respectively. We explore the parameter space for the lightest right-handed neutrino mass $M_{1}\in[10^{2},10^{14}]$ GeV and washout parameter $K$ that depends on the light-heavy neutrino Yukawa couplings $\lambda$, in the weak ($K<1$) and strong ($K>1$) washout regimes. The resulting novel features compatible with observed baryon asymmetry are detectable by future experiments like LISA and ET. By estimating signal-to-noise ratio (SNR) for upcoming GW experiments, we investigate the effect of the scalar mass $M_{\phi}$ and reheating temperature $T_{\phi}$, which depends on the $\phi-N$ Yukawa couplings $y_{N}$.
@@ I Introduction
The Standard Model (SM) of particle physics is unable to explain tiny neutrino masses and the observed matter-antimatter asymmetry of the Universe. Neutrino oscillation experiments involving solar Super-Kamiokande:2001bfk ; Super-Kamiokande:2002ujc ; SNO:2002tuh ; Super-Kamiokande:2005mbp ; Super-Kamiokande:2016yck ; Borexino:2015axw , atmospheric IceCube:2017lak ; ANTARES:2018rtf and reactor KamLAND:2008dgz ; T2K:2011ypd ; DoubleChooz:2011ymz ; T2K:2013ppw neutrinos provide evidence that neutrinos have mass and the flavor states mix due to the propagation of multiple mass eigenstates. This however only quantifies the squared mass differences of the neutrinos and not the absolute mass scales involved. On the other hand, the $\beta$-decay experiment KATRIN KATRIN:2021uub gives a stringent direct limit on the absolute value of neutrino mass scale, $m_{\nu}<0.8$ eV.
@@
which is in agreement with the values of abundances of light elements extracted from BBN (Big Bang Nucleosynthesis) data Fields:2019pfx .
@@
The discovery of Gravitational Waves (GWs) from black hole mergers by LIGO and Virgo collaboration LIGOScientific:2016aoc ; LIGOScientific:2016sjg and evidence for measurements of stochastic GW background by pulsar timing array (PTA) observations Carilli:2004nx ; Janssen:2014dka ; Weltman:2018zrl ; EPTA:2015qep ; EPTA:2015gke ; NANOGrav:2023gor ; NANOGrav:2023hvm have lead to several new physics cases to detect GWs of primordial origins as well. Particularly in the context of baryogenesis via leptogenesis, cosmological pathways to probe such high-scale physics involve predictions of the CMB spectral indices Ghoshal:2022fud or gravitational waves from local cosmic strings Dror:2019syi ; Saad:2022mzu ; DiBari:2023mwu ; Blasi:2020wpy , global cosmic strings Fu:2023nrn , domain walls Barman:2022yos ; King:2023cgv , nucleating and colliding vacuum bubbles Dasgupta:2022isg ; Borah:2022cdx , other topological defects Dunsky:2021tih , graviton bremmstrahlung Ghoshal:2022kqp , inflationary tensor perturbations propagating as GW Berbig:2023yyy ; Borboruah:2024eal , primordial blackholes Perez-Gonzalez:2020vnz ; Datta:2020bht ; JyotiDas:2021shi ; Barman:2021ost ; Bernal:2022pue ; Bhaumik:2022pil ; Bhaumik:2022zdd ; Ghoshal:2023fno , and exploring the oscillatory features of the primordial non-gaussianity and curvature bi-spectrum and tri-spectrum Cui:2021iie ; Fong:2023egk .
@@
Under this circumstance, we need more such probes of BSM physics and form a synergy between cosmological observations and laboratory experiments involving seesaw and low-scale leptogenesis. In this paper, we take a first step in this direction and investigate a general scenario where a heavy scalar denoted by $\phi$ may decay to create an initial non-thermal abundance of right-handed neutrinos (RHN) whose subsequent decays lead to leptogenesis satisfying Eq. (1). We will be concerned with Gravitational Waves that are generated due to quantum tensor fluctuations of the metric during inflation Grishchuk:1974ny ; Starobinsky:1979ty ; Rubakov:1982df ; Guzzetti:2016mkm . These GWs can act as a book-keeping of the expansion history of our Universe Seto:2003kc ; Boyle:2005se ; Boyle:2007zx ; Kuroyanagi:2008ye ; Nakayama:2009ce ; Kuroyanagi:2013ns ; Jinno:2013xqa ; Saikawa:2018rcs ; Chen:2024roo . Since the detailed time evolution of the Hubble rate during the expansion determines various PGW frequencies that are red-shifted to the present day, thus allowing us to investigate standard and non-standard thermal history of the universe in the pre-BBN era Bernal:2020ywq ; Nakayama:2008ip ; Nakayama:2008wy ; Kuroyanagi:2011fy ; Buchmuller:2013lra ; Buchmuller:2013dja ; Jinno:2014qka ; Kuroyanagi:2014qza ; Ghoshal:2023sfa ; Ghoshal:2022ruy ; Chen:2024roo . In this context, since we are interested in low scale seesaw and leptogenesis, we will study a scenario where $\phi$ dominates the energy budget of the Universe which leads to interesting GW spectral shapes. On one hand, $\phi$ decay will facilitate low-scale leptogenesis and on the other hand, it is also responsible for the generation of GW spectral shapes that can be measured by various GW detectors. In particular, we consider an epoch of intermediate matter domination by $\phi$ particles affecting the post-inflationary evolution of the Universe Scherrer:1984fd ; Kolb:1990vq ; Bezrukov:2009th .
@@
We will study GW probes of both thermal and non-thermal low scale leptogenesis scenarios. The leptogenesis scale is brought down mainly by the resonance effect that enhances the CP parameter $\epsilon$. Non-thermal leptogenesis is achieved via the scalar decay to RHNs. Non-thermal production of RHN creates an excess initial abundance of RHNs after the scalar has decayed, compared to the thermal case. Depending on reheating temperature $T_{\phi}$ and washout parameter $K$, the final efficiency factor $\kappa_{f}$ may get enhanced compared to the thermal case due to this excess initial abundance Hahn-Woernle:2008tsk ; Zhang:2023oyo . Unlike Berbig:2023yyy , here the RHNs need not dominate the Universe since the scalar is assumed to dominate the energy density of the Universe which distorts inflationary GWs. Therefore we do not need to stick to only the $K\ll 1$ condition of Berbig:2023yyy for RHN domination. Note that RHNs can still dominate after $\phi$ decays, depending on parameters.
@@
The decays occur very far away from thermal equilibrium, therefore it is bound to release a large amount of entropy dump, which in turn dilutes the energy density of primordial GWs that enter the horizon before the decay. We show that if the scalar couples directly to the radiation bath, under appropriate conditions RHNs produced non-thermally from $\phi$ thermalize quickly with the radiation bath, in which case entropy is injected in two stages, one from scalar decay and the other from RHN decay if they dominate the energy budget of the Universe before decaying. In this case, two knee-like features occur in the GW spectrum due to the scalar decay which we show for the first time. We will show the capability of several upcoming GW detectors like LISA, DECIGO, BBO, etc. in probing the parameter space for such leptogenesis, particularly in regions where it is impossible to test via laboratory or astrophysical experiments.
@@
The paper is organized as follows: In Sec. II, we give a brief overview of the thermal leptogenesis mechanism and the Davidson-Ibarra bound on heavy neutrino masses. We consider resonant leptogenesis to bring the leptogenesis scale down. In Sec. III, we discuss the non-thermal leptogenesis scenario from scalar decay and the relevant Boltzmann equations. In Sec. IV, we identify three classes of scenarios relevant to GW observations, two of which correspond to non-thermal and one to thermal leptogenesis. We provide analytical approximations of matter-radiation equilibrium temperature and dilution factor from entropy injection for these classifications. Then in Sec. V we discuss primordial gravitational waves from inflation with and without an intermediate matter domination epoch. We also introduce a transfer function to encapsulate a 2-step entropy injection. This section is followed by our results in Sec. VI. Finally, we conclude this paper in Sec. VII.
II Thermal Leptogenesis
@@
Before discussing the effect of scalar field reheating on leptogenesis let us first we give a brief review of standard thermal leptogenesis in this section. Initially proposed by Fukugita and Yanagida in 1986 Fukugita:1986hr ; Yanagida:2005hz , thermal leptogenesis hinges on the out-of-equilibrium decay of RHNs, which are thermally produced through scattering processes in the early Universe from the thermal bath of the primordial SM plasma. The Lagrangian is given by,
[…]
@@
where $\mathcal{L}_{\rm SM}$ is the SM Lagrangian, $N$ represents RHNs, $\tilde{H}=i\sigma_{2}H^{*}$ where $H$ is the SM Higgs doublet and $L$ is the SM leptonic doublet. Here $M_{N}=\text{diag}[M_{1},M_{2},M_{3}]$ is the RHN Majorana mass matrix and $\lambda$ encompasses the light-heavy Yukawa couplings. Due to the Majorana mass term of RHNs, after electroweak symmetry breaking (EWSB), the familiar Type-I seesaw mechanism Yanagida:1979as provides mass to the light neutrinos, $m_{\nu}=v^{2}\lambda M_{N}^{-1}\lambda^{T}$, where $v\sim 174$ GeV is the vacuum expectation value (VEV) of the SM Higgs. On the other hand, RHNs of Majorana nature, can decay into both leptons and anti-leptons through the Yukawa coupling matrix $\lambda$, thereby violating lepton number (LNV),
@@
is the total decay width of $N_{1}$ in the rest frame of the neutrino. The lepton asymmetry can be re-written as $B-L$ asymmetry denoted by the number density $n_{B-L}$, which is subsequently converted into the baryon asymmetry through $B+L$ conserving electroweak sphaleron processes Harvey:1990qw ; Khlebnikov:1988sr ; Luty:1992un . When there is no initial $B-L$ asymmetry, the final baryon asymmetry, denoted as baryon-to-photon number density ratio $\eta_{B}$, is expressed as Buchmuller:2004nz ,
$\eta_{B}\equiv\frac{n_{B}-n_{\overline{B}}}{n_{\gamma}}=-\frac{3}{4}\frac{a_{\text{sph}}}{f}\,\epsilon_{1}\kappa_{f}\equiv-d\,\epsilon_{1}\kappa_{f}\simeq 0.96\times 10^{-2}\,\epsilon_{1}\kappa_{f}$ (6)
where $a_{\text{sph}}=28/79$ Arnold:1987mh ; Khlebnikov:1988sr ; Harvey:1990qw is the efficiency of the sphaleron processes of converting $B-L$ asymmetry into baryon asymmetry and $f=2387/86$ is the dilution factor due to production of photons since the onset of leptogenesis till recombination assuming an isentropic expansion of the Universe. And $\kappa_{f}$ is the efficiency parameter Buchmuller:2004nz , which is independent of $\epsilon_{1}$ and which parameterizes the efficiency of lepton asymmetry production depending on the initial $N_{1}$ abundance and washout. It can be calculated by solving the relevant Boltzmann equations governing the evolution of the $N_{1}$ abundance and $B-L$ asymmetry Nardi:2007jp ; Basboll:2006yx . Its final value is normalized to unity if thermal initial abundance of $N_{1}$ is assumed and there is no washout.
@@ II.1 Conditions for Departure from Thermal Equilibrium
where $\mathcal{H}(T\sim M_{i})\equiv\mathcal{H}(M_{i})=\sqrt{\frac{8\pi^{3}g_{\ast}}{90}}\frac{M_{i}^{2}}{M_{\text{pl}}}$ is the Hubble parameter with $M_{\text{pl}}=1.22\times 10^{19}$ GeV and $g_{*}$ is the relativistic degrees of freedom which is approximately equal to $106.75$ at the high temperatures under discussion. It can be shown that the above condition is satisfied when the effective neutrino mass, defined by,
$\tilde{m}_{1}=\frac{(m_{D}^{\dagger}m_{D})_{11}}{M_{1}}$ (8)
@@
The ratio is called the decay parameter,
@@
which determines whether $N_{1}$ decays are in thermal equilibrium or not. The case $K<1$ is called the weak washout regime, where Yukawa interactions of $N_{1}$ with light neutrinos via the SM Higgs are weak, hence washout of any initial asymmetry via reverse processes is small. Therefore leptogenesis is unavoidable in this regime. However in the so called strong washout regime ($K>1$), although Yukawa interactions are strong, leptogenesis can still occur and the final asymmetry depends on the interplay between decay, inverse decay and scattering of RHNs Buchmuller:2004nz .
It turns out the final efficiency parameter $\kappa_{f}$ can be estimated analytically for both weak and strong washout regime Buchmuller:2004nz . In the strong washout regime, $\kappa_{f}$ is estimated by,
@@
where $\text{where }z_{B}(K)$ $\simeq 1+\frac{1}{2}\text{ ln}\left(1+\frac{\pi K^{2}}{1024}\left[\text{ ln}\left(\frac{3125\pi K^{2}}{1024}\right)\right]^{5}\right)$ (14)

### v2

@@ Abstract
We study non-thermal resonant leptogenesis in a general setting where a heavy majoron $\phi$ decays to right-handed neutrinos (RHNs) whose further out-of-equilibrium decay generates the required lepton asymmetry. Domination of the energy budget of the Universe by the $\phi$ or the RHNs alters the evolution history of the primordial gravitational waves (PGW) of inflationary origin, which re-enter the horizon after inflation, modifying the spectral shape. The decays of $\phi$ and RHNs release entropy into the early Universe while nearly degenerate RHNs facilitate low and intermediate-scale leptogenesis. A characteristic damping of the GW spectrum resulting in knee-like features would provide evidence for low-scale non-thermal leptogenesis. We explore the parameter space for the lightest right-handed neutrino mass $M_{1}\in[10^{2},10^{14}]$ GeV and washout parameter $K$ that depends on the light-heavy neutrino Yukawa couplings $\lambda$, in the weak ($K<1$) and strong ($K>1$) washout regimes. The resulting novel features compatible with observed baryon asymmetry are detectable by future experiments like LISA and ET. By estimating signal-to-noise ratio (SNR) for upcoming GW experiments, we investigate the effect of the majoron mass $M_{\phi}$ and reheating temperature $T_{\phi}$, which depends on the $\phi-N$ Yukawa couplings $y_{N}$.
@@ I Introduction
The Standard Model (SM) of particle physics cannot explain tiny neutrino masses and the observed matter-antimatter asymmetry of the Universe. Neutrino oscillation experiments involving solar Super-Kamiokande:2001bfk ; Super-Kamiokande:2002ujc ; SNO:2002tuh ; Super-Kamiokande:2005mbp ; Super-Kamiokande:2016yck ; Borexino:2015axw , atmospheric IceCube:2017lak ; ANTARES:2018rtf and reactor KamLAND:2008dgz ; T2K:2011ypd ; DoubleChooz:2011ymz ; T2K:2013ppw neutrinos provide evidence that neutrinos have mass and the flavor states mix due to the propagation of multiple mass eigenstates. This however only quantifies the squared mass differences of the neutrinos and not the absolute mass scales involved. On the other hand, the $\beta$-decay experiment KATRIN KATRIN:2021uub gives a stringent direct limit on the absolute value of neutrino mass scale, $m_{\nu}<0.8$ eV.
@@
This agrees with the values of abundances of light elements extracted from BBN (Big Bang Nucleosynthesis) data Fields:2019pfx .
@@
The discovery of Gravitational Waves (GWs) from black hole mergers by LIGO and Virgo collaboration LIGOScientific:2016aoc ; LIGOScientific:2016sjg and evidence for measurements of stochastic GW background by pulsar timing array (PTA) observations Carilli:2004nx ; Janssen:2014dka ; Weltman:2018zrl ; EPTA:2015qep ; EPTA:2015gke ; NANOGrav:2023gor ; NANOGrav:2023hvm have lead to several new physics cases to detect GWs of primordial origins as well. Particularly in the context of baryogenesis via leptogenesis, cosmological pathways to probe such high-scale physics involve predictions of the CMB spectral indices Ghoshal:2022fud or gravitational waves from local cosmic strings Dror:2019syi ; Saad:2022mzu ; DiBari:2023mwu ; Blasi:2020wpy , global cosmic strings Fu:2023nrn , domain walls Barman:2022yos ; King:2023cgv , nucleating and colliding vacuum bubbles Dasgupta:2022isg ; Borah:2022cdx , other topological defects Dunsky:2021tih , graviton bremmstrahlung Ghoshal:2022kqp , gravitational production of neutrinos Haque:2023zhb , inflationary tensor perturbations propagating as GW Berbig:2023yyy ; Borboruah:2024eal , primordial blackholes Perez-Gonzalez:2020vnz ; Datta:2020bht ; Barman:2024slw ; JyotiDas:2021shi ; Barman:2021ost ; Bernal:2022pue ; Bhaumik:2022pil ; Bhaumik:2022zdd ; Ghoshal:2023fno , and exploring the oscillatory features of the primordial non-gaussianity and curvature bi-spectrum and tri-spectrum Cui:2021iie ; Fong:2023egk .
@@
Under this circumstance, we need more such probes of BSM physics and form a synergy between cosmological observations and laboratory experiments involving seesaw and low-scale leptogenesis. In this paper, we take a first step in this direction and investigate a general scenario where a heavy scalar called majoron, denoted by $\phi$, may decay to create an initial non-thermal abundance of right-handed neutrinos (RHNs) whose subsequent decay leads to leptogenesis satisfying Eq. (1). This majoron $\phi$ is different from inflaton. Inflaton itself is not the direct concern of this work although we use the inflationary reheating temperature $T_{\rm RH}$ in our calculations. We will be concerned with gravitational waves that are generated due to quantum tensor fluctuations of the metric during inflation Grishchuk:1974ny ; Starobinsky:1979ty ; Rubakov:1982df ; Haque:2021dha ; Guzzetti:2016mkm . These GWs can act as a book-keeping of the expansion history of our Universe Seto:2003kc ; Boyle:2005se ; Boyle:2007zx ; Kuroyanagi:2008ye ; Nakayama:2009ce ; Kuroyanagi:2013ns ; Jinno:2013xqa ; Saikawa:2018rcs ; Haque:2023zhb ; Chen:2024roo . Since the detailed time evolution of the Hubble rate during the expansion determines various PGW frequencies that are red-shifted to the present day, thus allowing us to investigate standard and non-standard thermal history of the universe in the pre-BBN era Bernal:2020ywq ; Nakayama:2008ip ; Nakayama:2008wy ; Kuroyanagi:2011fy ; Buchmuller:2013lra ; Buchmuller:2013dja ; Jinno:2014qka ; Kuroyanagi:2014qza ; Haque:2021dha ; Ghoshal:2023sfa ; Ghoshal:2022ruy ; Chen:2024roo . In this context, since we are interested in low-scale seesaw and leptogenesis, we will study a scenario where $\phi$ dominates the energy budget of the Universe which leads to interesting GW spectral shapes. On one hand, $\phi$ decay will facilitate low-scale leptogenesis and on the other hand, it is also responsible for the generation of GW spectral shapes that can be measured by various GW detectors. In particular, we consider an epoch of intermediate matter domination by $\phi$ particles affecting the post-inflationary evolution of the Universe Scherrer:1984fd ; Barman:2024slw ; Kolb:1990vq ; Bezrukov:2009th .
@@
In this model, non-thermal leptogenesis is achieved via the majoron decay to RHNs. The leptogenesis scale is brought down mainly by the resonance effect that enhances the CP parameter $\epsilon$. RHNs are produced in an excess initial abundance after the majoron has decayed, compared to the thermal case. Depending on reheating temperature $T_{\phi}$ and washout parameter $K$, the final efficiency factor $\kappa_{f}$ may get enhanced compared to the thermal case due to this excess initial abundance Hahn-Woernle:2008tsk ; Zhang:2023oyo . Unlike Berbig:2023yyy , here the RHNs need not dominate the Universe since the majoron is assumed to dominate the energy density of the Universe which distorts inflationary GWs. Therefore we do not need to stick to only the $K\ll 1$ condition of Berbig:2023yyy for RHN domination. Note that RHNs can still dominate after $\phi$ decays, depending on parameters.
@@
The decays occur far away from thermal equilibrium, therefore it is bound to release a large amount of entropy dump, which in turn dilutes the energy density of primordial GWs that enter the horizon before the decay. We show that a 2-step entropy injection scenario is possible if the RHNs produced are highly relativistic at production, which eventually become non-relativistic and then decay at a lower temperature. In this case, two knee-like features occur in the GW spectrum due to the majoron and RHN decays which we show for the first time. We will show the capability of several upcoming GW detectors like LISA, DECIGO, BBO, etc. in probing the parameter space for such leptogenesis, particularly in regions where it is impossible to test via laboratory or astrophysical experiments.
In this work a framework of non-thermal resonant leptogenesis from the decay chain $\phi\to N_{1,2}\to H+L$, includes entropy release from both $\phi$ and RHNs, spans RHN masses $10^{2}$$10^{14}$ GeV and washout strengths $K\sim 10^{-20}$$10^{4}$, systematically classifies scenarios, and highlights knee-like damping and kinked GW spectra with SNR estimates across detectors. Unlike the above studies, we focus on a nearly-degenerate RHN sector decoupled from the thermal bath, providing a general yet distinctive link between majoron-induced resonant leptogenesis and observable GW features [^1] See Refs. DEramo:2019tit ; Datta:2022tab ; Chianese:2024nyw ; Samanta:2025jec for ideas related to probing non-standard expansion histories and leptogenesis through inflationary gravitational waves..
@@
The paper is organized as follows: In Sec. II, we give a brief review of the thermal leptogenesis mechanism and the Davidson-Ibarra bound on heavy neutrino masses. We consider resonant leptogenesis to bring the leptogenesis scale down. In Sec. III, we discuss the majoron model that facilitates non-thermal leptogenesis in details and the dynamics of majoron and RHN decays. We construct the relevant Boltzmann equations. In Sec. IV, we identify three classes of scenarios relevant to GW observations. We provide analytical approximations of matter-radiation equilibrium temperature and dilution factor from entropy injection for these classifications. Then in Sec. V we discuss primordial gravitational waves from inflation with and without an intermediate matter domination epoch. We also introduce a transfer function to encapsulate a 2-step entropy injection. This section is followed by our results in Sec. VI. Finally, we conclude this paper in Sec. VII. In Appendix A, we explicitly demonstrate a scenario where washout factor can be tiny while keeping the CP asymmetry large. In Appendix B we describe the Boltzmann equations in terms of comoving quantities.
II Review of Thermal Leptogenesis
@@
Before discussing the effect of majoron field reheating on leptogenesis, let us first give a brief review of standard thermal leptogenesis. Initially proposed by Fukugita and Yanagida in 1986 Fukugita:1986hr ; Yanagida:2005hz , thermal leptogenesis hinges on the out-of-equilibrium decay of RHNs, which are thermally produced through scattering processes in the early Universe from the thermal bath of the primordial SM plasma. The Lagrangian is given by,
$\mathcal{L}_{\rm thermal}={\cal L}_{\rm SM}+i\overline{N}\not{\partial}N-\left(\lambda\overline{L}\tilde{H}N+\frac{M_{N}}{2}\overline{N^{C}}N+\text{h.c}\right)$ (2)
@@
where $\mathcal{L}_{\rm SM}$ is the SM Lagrangian, $N$ represents RHNs, $\tilde{H}=i\sigma_{2}H^{*}$ where $H$ is the SM Higgs doublet and $L$ is the SM leptonic doublet. Here $M_{N}=\text{diag}[M_{1},M_{2},M_{3}]$ is the RHN Majorana mass matrix and $\lambda$ encompasses the light-heavy Yukawa couplings. Due to the Majorana mass term of RHNs, after electroweak symmetry breaking (EWSB), the familiar Type-I seesaw mechanism Yanagida:1979as provides mass to the light neutrinos, $m_{\nu}=v^{2}\lambda^{T}M_{N}^{-1}\lambda$, where $v\sim 174$ GeV is the vacuum expectation value (VEV) of the SM Higgs. On the other hand, RHNs of Majorana nature can decay into both leptons and anti-leptons through the Yukawa coupling matrix $\lambda$, thereby violating lepton number (LNV),
@@
is the total decay width of $N_{1}$ in the rest frame of the neutrino. The lepton asymmetry can be re-written as BLB-L asymmetry denoted by the number density $n_{B-L}$, which is subsequently converted into the baryon asymmetry through $B+L$ conserving electroweak sphaleron processes Harvey:1990qw ; Khlebnikov:1988sr ; Luty:1992un . When there is no initial BLB-L asymmetry, the final baryon asymmetry, denoted as baryon-to-photon number density ratio $\eta_{B}$, is expressed as Buchmuller:2004nz ,
$\eta_{B}\equiv\frac{n_{B}-n_{\overline{B}}}{n_{\gamma}}=\frac{3}{4}\frac{a_{\text{sph}}}{f}\,\epsilon_{1}\kappa_{f}\simeq 0.96\times 10^{-2}\,\epsilon_{1}\kappa_{f}$ (6)
where $a_{\text{sph}}=28/79$ Arnold:1987mh ; Khlebnikov:1988sr ; Harvey:1990qw is the efficiency of the sphaleron processes of converting BLB-L asymmetry into baryon asymmetry and $f=2387/86$ is the dilution factor due to production of photons since the onset of leptogenesis till recombination assuming an isentropic expansion of the Universe. And $\kappa_{f}$ is the efficiency parameter Buchmuller:2004nz , which is independent of $\epsilon_{1}$ and which parameterizes the efficiency of lepton asymmetry production depending on the initial $N_{1}$ abundance and washout. It can be calculated by solving the relevant Boltzmann equations governing the evolution of the $N_{1}$ abundance and BLB-L asymmetry Nardi:2007jp ; Basboll:2006yx . Its final value is normalized to unity if a thermal initial abundance of $N_{1}$ is assumed and there is no washout.
@@ II.1 Conditions for Departure from Thermal Equilibrium
where $\mathcal{H}(T\sim M_{1})\equiv\mathcal{H}(M_{1})=\sqrt{\frac{8\pi^{3}g_{\ast}}{90}}\frac{M_{1}^{2}}{M_{\text{pl}}}$ is the Hubble parameter with $M_{\text{pl}}=1.22\times 10^{19}$ GeV and $g_{*}$ is the relativistic degrees of freedom which is approximately equal to $106.75$ at the high temperatures under discussion. It can be shown that the above condition is satisfied when the effective neutrino mass, defined by,
$\tilde{m}_{1}=\frac{(m_{D}m_{D}^{\dagger})_{11}}{M_{1}}$ (8)
@@
The ratio is called the decay parameter or the washout parameter,
@@
which determines whether $N_{1}$ decays are in thermal equilibrium or not. The case $K<1$ is called the weak washout regime, where Yukawa interactions of $N_{1}$ with light neutrinos via the SM Higgs are weak, hence washout of any initial asymmetry via reverse processes is small. Therefore leptogenesis is unavoidable in this regime. However in the so-called strong washout regime ($K>1$), although Yukawa interactions are strong, leptogenesis can still occur and the final asymmetry depends on the interplay between decay, inverse decay, and scattering of RHNs Buchmuller:2004nz .
Within type-I seesaw framework, the Yukawa couplings $\lambda$ can be expressed by the Casas-Ibarra parametrization Casas:2001sr ,
$\lambda=\frac{1}{v}\,M_{N}^{1/2}\,R\,m_{\nu}^{1/2}U_{\rm PMNS}^{\dagger},$ (11)
where $m_{\nu}=\text{diag}(m_{1},m_{2},m_{3})$, $M_{N}=\text{diag}(M_{1},M_{2},M_{3})$ and $R$ is a complex orthogonal matrix. The effective neutrino masses,
$\tilde{m}_{i}=\frac{\left|\lambda\lambda^{\dagger}\right|_{ii}v^{2}}{M_{i}}=\sum_{j}m_{j}|R_{ij}|^{2}$ (12)
correspond to the physical neutrino masses $m_{i}$ if $R$ is diagonal. This means, assuming normal ordering $m_{1}<m_{2}<m_{3}$ and $m_{1}=0$, for small off-diagonal elements of $R$ matrix, the effective neutrino mass $\tilde{m}_{1}\rightarrow 0$, hence the washout parameter $K\rightarrow 0$. However, in this case, $\tilde{m}_{2}\rightarrow 10^{-3}$ eV since $m_{2}=\sqrt{m_{1}^{2}+\Delta m_{\rm atm}^{2}}\sim 8.6\times 10^{-3}$ eV, implying the washout factor $K_{2}=\tilde{m}_{2}/m_{*}$ for $N_{2}$ remains $\mathcal{O}(1)$. Similarly $K_{3}$ also remains large. It can be shown that, in general, $\tilde{m}_{1}$ is larger than the lightest active neutrino mass Fujii:2002jw ,
$\tilde{m}_{1}>\text{Min}[m_{1},m_{2},m_{3}]$ (13)
It turns out the final efficiency parameter $\kappa_{f}$ can be estimated analytically for both weak and strong washout regimes Buchmuller:2004nz . In the strong washout regime, $\kappa_{f}$ is estimated by,
@@
$\text{where }z_{B}(K)$ $\simeq 1+\frac{1}{2}\text{ ln}\left(1+\frac{\pi K^{2}}{1024}\left[\text{ ln}\left(\frac{3125\pi K^{2}}{1024}\right)\right]^{5}\right)$ (17)